# HyperSIGMA: Hyperspectral Intelligence Comprehension Foundation Model

### v1

@@ 4 Methodology
The construction of HyperSIGMA involves three main steps: initializing model weights through pre-training, enhancing model structure with SSA, and fusing spatial-spectral features. Following typical HSI processing strategies [39], we use two parallel subnetworks to extract spatial and spectral features. First, inspired by [164, 84], we obtain the model weights using MAE [146] pre-training on HyperGlobal-450K. Importantly, the spatial and spectral networks are pre-trained separately. Next, SSA is integrated into the model to enhance its structure. Lastly, we fuse the spatial and spectral information to enhance the representation of the extracted features, resulting in the final HyperSIGMA model. Technique details will be presented in the following sections.
@@ 4.1.2 Pre-training of Spatial Subnetwork
Similar to many multispectral FMs [97, 3, 102, 106], we use ViT [76] as the backbone for the spatial subnetwork. Following successful practices [164, 84, 97, 88, 94, 95, 85, 102, 3], we employ MAE [146] for pre-training. The only modification from the original implementation on natural images is adjusting the input channel of the patch embedding layers to match the number of channels in the input HSIs.

### v2

@@ 4 Methodology
The construction of HyperSIGMA involves three main steps: initializing model weights through pre-training, enhancing model structure with SSA, and fusing spatial-spectral features. Following typical HSI processing strategies [19], we use two parallel subnetworks to extract spatial and spectral features. First, inspired by [49, 31], we obtain the model weights using MAE [46] pre-training on HyperGlobal-450K. Importantly, the spatial and spectral networks are pre-trained separately. Next, SSA is integrated into the model to enhance its structure. Lastly, we fuse the spatial and spectral information to enhance the representation of the extracted features, resulting in the final HyperSIGMA model. Technical details will be presented in the following sections.
@@ 4.1.2 Pre-training of Spatial Subnetwork
Similar to many multispectral foundation models [42, 3], we use ViT [29] as the backbone for the spatial subnetwork. Following successful practices [49, 31, 42, 3], we employ MAE [46] for pre-training. The only modification from the original implementation on natural images is adjusting the input channel of the patch embedding layers to match the number of channels in the input HSIs.